# Video[Video][video]

### v1

@@ 2 Background and prior work
Many works describe microscopic image formation for individual contrast modes including fluorescence contrast [13, 14], phase and absorption contrast [15, 16], and polarization-resolved birefringence contrast [17, 18]. Recently, multiple papers have developed computational imaging methods for integrative measurements of specimen properties, but with models specific to label-free contrast [3, 19, 20, 21] or fluorescence contrast [22, 23]. The waveOrder framework unifies and extends these models to multi-contrast and multi-channel imaging systems that include scalar and polarization-resolved imaging with or without fluorescence labels.
@@
We start by describing a framework for reconstructing material properties from microscopic imaging data. By representing material properties and data as vectors, we describe our reconstructions as the solution to an optimization problem. Using broadly applicable assumptions of linearity, shift invariance, and weak scattering, we describe all contrasts as the result of banks of point spread functions. We describe the physics behind the major contrast mechanisms and compute their transfer functions from illumination, scattering, and detection models. Throughout, we use carefully chosen basis functions to enable the physical interpretation of light and material properties. We describe all notation as we introduce it, and we collect all symbols in Section 10.
@@ 3.1 Objects and data as vectors, imaging as an operator
Figure 1: A framework for volumetric microscopy of phase, absorption, birefringence, diattenuation, fluorescence density, and fluorescence orientation. (a) Microscopes encode (i) true material properties into (ii) multi-contrast multi-channel volumetric datasets by changing spectral and polarization illumination and detection filters. Physics-informed models of the encoding process, $\mathcal{H}$, lead to physics-informed reconstruction operators, $\mathcal{R}$, that enable (iii) reconstructions of underlying material properties. (iv) We find that physics-informed reconstructions are useful for improving performance on downstream observation, biophysical, and learned tasks. (b) We model the encoding process of an (i) off-the-shelf inverted microscope stand outfitted with (ii) an incoherent source, variable spectral and polarization filters, and a polarization-sensitive detector. Two classes of contrast are available: label-free contrast where (iii) direct fields scatter coherently from the sample and direct-scatter interference generates contrast, and (iv) fluorescence contrast where direct fields that excite inelastic-scattering fluorophores are filtered so that scatter-scatter interference generates contrast. We model coherent scatterers as (v) a volume of scattering potential tensors that model the relationship between incident electric fields and induced dipoles moments. We model inelastic fluorescence scatterers as (vi) a volume of dipole emission moments. Both coherent and incoherent dipoles can be linear, circular, and elliptical, though fluorescent dipoles used in biological microscopy are primarily linear. (c) Expanding the scattering potential tensor onto spherical harmonic tensors enables interpretation in terms of phase, absorption, birefringence, and diattenuation. (d) Expanding dipole emitters into Cartesian components with complex coefficients enables interpretation of single emitters’ orientations. (e) Expanding measurable data into Stokes parameters enables modeling of all possible polarization measurements. See also Video 19.
@@ 3.2 Reconstructing object properties
After choosing an objective function, the minimization problem needs to be solved, an often challenging task. Fortunately, if $\mathcal{H}$ is linear then the Tikhonov-regularized least-squares objective can be minimized in a single step described in Supplement 14, generating efficient noise-tolerant estimates of material properties, Figure 1a, iii.
@@ 3.4 Physically interpretable basis functions
When we illuminate a label-free sample, the 3D induced dipole moment is the product of the incident field and a 3 $\times$ 3 matrix called the permittivity tensor [28, 21]. By convention, we change to a unitless quantity and subtract the isotropic background (Supplement 15) to arrive at a complete set of label-free sample properties—the complex-valued 3 $\times$ 3 matrix called the scattering potential tensor, $f^{(\text{lf})}_{ij}$. Each entry of the scattering potential tensor can be interpreted directly (e.g. the complex-valued $f^{(\text{lf})}_{xz}$ is the relative magnitude and phase of the $x$ component of the dipole induced by a $z$-oriented field), but this interpretation can be challenging to understand physically. To improve physical interpretability, we expand the scattering potential tensor onto the spherical harmonic tensors, a set of nine 3 $\times$ 3 matrices whose complex-valued expansion coefficients can be directly interpreted in terms of phase, absorption, birefringence, and diattenuation. We schematize each of these spherical harmonic tensors in Figure 1c by drawing each tensor’s eigenvalues and eigenvectors, and we describe the spherical harmonic tensor basis in detail in Supplement 15.
@@
Finally, we express our data in terms of the Stokes parameters, a set of four real-valued parameters that are physically interpretable as the intensities measured behind various polarizing filters. The Stokes parameters can also be interpreted as the coefficients of the electric field’s second-moment matrix expanded onto the Pauli matrices (Figure 1e, Supplement 16).
@@ 3.5 Linear contrast-separable shift-invariant imaging systems
Under approximations that are applicable to a wide range of microscopes (details in Section 5 and Supplement 17), we can model our imaging system by splitting our material properties, data channels, and forward model into independent groups called contrast modes, e.g. one label-free mode and several fluorescence modes. Each of these contrast modes is approximately linear across object properties, and each property-to-channel mapping is approximately spatially linear and shift invariant. Therefore, we can express our imaging model as
@@ 3.6 Summary of transfer functions
Figure 2: Key components of the forward model. (a) A dipole emitter (blue arrow) emits polarized fields (red arrows) in an anisotropic pattern (black line, radius proportional to power). (b) Arbitrary linear, circular, and elliptical dipole emissions can be modeled with the Green’s tensor spectrum $G_{ij}(\boldsymbol{\nu})$. Each surface shows the quarter-maximum intensity of on-shell spectral field components (rows) $e_{x}$, $e_{y}$, and $e_{z}$ emitted by dipoles oriented along (columns) $\hat{\textbf{x}}$, $\hat{\textbf{y}}$, and $\hat{\textbf{z}}$, with relative phase shown in color (see color rose). Orthogonal slices through the Green’s tensor spectrum and the Green’s tensor are shown in Supplement Figure 8. (c–f) 3D support of various transfer functions with phase encoded in color. (c-d) Real label-free vector transfer functions with $\text{NA}_{\text{det}}=0.75$ and circularly polarized illumination that expresses polarization-resolved data (rows) as outputs of filters that modulate material properties (columns) with (c) low-NA illumination $\text{NA}_{\text{ill}}=0.1$ and (d) high-NA illumination $\text{NA}_{\text{ill}}=0.5$. (e) Scalar label-free transfer function with $\text{NA}_{\text{ill}}=0.5$ and $\text{NA}_{\text{det}}=0.75$. (f) Scalar fluorescence transfer function with $\text{NA}_{\text{det}}=0.75$. See also Video 19.
All transfer functions can be expressed as products and autocorrelations of the illumination, scattering, and detection models, see Table 1 and Supplement 17. We refer to the complete transfer functions as vector models because they account for the complete vectorial nature of light and dipole scattering. We also include scalar models that ignore vector effects, which are reasonable approximations when unpolarized illumination and unpolarized detection are used on isotropic samples.
@@ 4.1 Label-free simulated and real test specimen
Figure 3: Demonstration of vectorial multi-channel reconstruction with experiment and simulation. We imaged (a) (i) a laser-etched spoke pattern of transverse radially anisotropic bubbles. (ii) We made volumetric measurements of the Stokes parameters (columns), and applied a multi-channel label-free reconstruction to (iii) recover material properties. (b) We (i) simulated the phantom’s material properties, (ii) simulated the imaging process, then (iii) simulated the reconstruction. (c) We encoded transverse birefringence properties into color, where brightness indicates the strength of the anisotropy and hue indicate the slow-axis orientation. We compared the orientation and phase reconstructions (rows) in ray- and wave-optics reconstructions (columns). (d) For each data and material property, we measured a series of azimuthal profiles at different radii on the spoke pattern and used the 10th–90th percentile modulation as an empirical estimate of the transverse modulation transfer function (MTF). (e) Similarly, we measured axial profiles through each data and material property. See also Video 19.
@@ 4.2 Cells and tissues
Figure 4: Reconstructions improve visualization and interpretation across length scales in (a) A549 cells, (b) zebrafish, and (c–e) zebrafish regions of interest. (i) Orientation data (left) and reconstructions (right). (ii) Brightfield data (left) and phase reconstructions (right). (iii) Fluorescence data (left) and reconstructions (right). See also Videos 19 and 19.
@@ 4.3 Choice of regularization parameter
waveOrder’s physics-informed reconstructions rely on a list of fixed parameters that specify imaging conditions (e.g., numerical apertures, wavelength, voxel sizes) and a single regularization parameter $\eta$, which penalizes large solutions and suppresses noise. Though principled ways to choose regularization parameters exist [34, 35], in practice, it is common to inspect a regularization sweep and choose an empirical solution (for example, Figure 5 and Supplement Figure 6). Regularization parameters that are too small result in noisy reconstructions, while regularization parameters that are too large result in blurry reconstructions. Additionally, we find that the transverse mid-band power, the power in transverse spatial frequencies between one-eighth and one-quarter of the cutoff, is a useful scalar metric that can indicate high-quality reconstructions—see the plateau in Figure 5(d).
@@ 5 Limitations
  • shift invariance, which excludes shift-variant point spread functions,
  • weak, single scattering, which excludes thick multiply scattering samples, e.g. older zebrafish and thick tissue slices,
@@
  • contrast-separable imaging, which excludes crosstalk between label-free and fluorescence contrast modes.
@@ 8 Acknowledgements
We thank Shashvat Mehta for developing the animation shown in Video 19. We thank the Scientific Computing Platform at CZ Biohub for enabling high-performance reconstructions.

### v2

@@ 2 Background and prior work
Many works describe microscopic image formation for individual contrast modes including fluorescence contrast [13, 14], phase and absorption contrast [15, 16], and polarization-resolved birefringence contrast [17, 18]. Recently, multiple papers have developed computational imaging methods for integrative measurements of specimen properties, but with models specific to label-free contrast [3, 19, 20, 21, 22] or fluorescence contrast [23, 24]. The waveOrder framework unifies and extends these models to multi-contrast and multi-channel imaging systems that include scalar and polarization-resolved imaging with or without fluorescence labels.
@@
We start by describing a framework for reconstructing material properties from microscopic imaging data. By representing material properties and data as vectors, we describe our reconstructions as the solution to an optimization problem. Using broadly applicable assumptions of linearity, shift invariance, and weak scattering, we describe all contrasts as the result of banks of point spread functions. We describe the physics behind the major contrast mechanisms and compute their transfer functions from illumination, scattering, and detection models. Throughout, we use carefully chosen basis functions to enable the physical interpretation of light and material properties. We describe all notation as we introduce it, and we collect all symbols in Supplement 1.
@@ 3.1 Objects and data as vectors, imaging as an operator
Figure 1: A framework for volumetric microscopy of phase, absorption, birefringence, diattenuation, fluorescence density, and fluorescence orientation. (a) Microscopes encode (i) true material properties into (ii) multi-contrast multi-channel volumetric datasets by changing spectral and polarization illumination and detection filters. Physics-informed models of the encoding process, $\mathcal{H}$, lead to physics-informed reconstruction operators, $\mathcal{R}$, that enable (iii) reconstructions of underlying material properties. (iv) We find that physics-informed reconstructions are useful for improving performance on downstream observation, biophysical, and learned tasks. (b) We model the encoding process of an (i) off-the-shelf inverted microscope stand outfitted with (ii) an incoherent source, variable spectral and polarization filters, and a polarization-sensitive detector. Two classes of contrast are available: label-free contrast where (iii) direct fields scatter coherently from the sample and direct-scatter interference generates contrast, and (iv) fluorescence contrast where direct fields that excite inelastic-scattering fluorophores are filtered so that scatter-scatter interference generates contrast. We model coherent scatterers as (v) a volume of scattering potential tensors that model the relationship between incident electric fields and induced dipoles moments. We model inelastic fluorescence scatterers as (vi) a volume of dipole emission moments. Both coherent and incoherent dipoles can be linear, circular, and elliptical, though fluorescent dipoles used in biological microscopy are primarily linear. (c) Expanding the scattering potential tensor onto spherical harmonic tensors enables interpretation in terms of phase, absorption, birefringence, and diattenuation. (d) Expanding dipole emitters into Cartesian components with complex coefficients enables interpretation of single emitters’ orientations. (e) Expanding measurable data into Stokes parameters enables modeling of all possible polarization measurements. See also Video 1.
@@ 3.2 Reconstructing object properties
After choosing an objective function, the minimization problem needs to be solved, an often challenging task. Fortunately, if $\mathcal{H}$ is linear then the Tikhonov-regularized least-squares objective can be minimized in a single step described in Supplement 5, generating efficient noise-tolerant estimates of material properties, Figure 1a, iii.
@@ 3.4 Physically interpretable basis functions
When we illuminate a label-free sample, the 3D induced dipole moment is the product of the incident field and a 3 $\times$ 3 matrix called the permittivity tensor [29, 21]. By convention, we change to a unitless quantity and subtract the isotropic background (Supplement 6) to arrive at a complete set of label-free sample properties—the complex-valued 3 $\times$ 3 matrix called the scattering potential tensor, $f^{(\text{lf})}_{ij}$. Each entry of the scattering potential tensor can be interpreted directly (e.g. the complex-valued $f^{(\text{lf})}_{xz}$ is the relative magnitude and phase of the $x$ component of the dipole induced by a $z$-oriented field), but this interpretation can be challenging to understand physically. To improve physical interpretability, we expand the scattering potential tensor onto the spherical harmonic tensors, a set of nine 3 $\times$ 3 matrices whose complex-valued expansion coefficients can be directly interpreted in terms of phase, absorption, birefringence, and diattenuation. We schematize each of these spherical harmonic tensors in Figure 1c by drawing each tensor’s eigenvalues and eigenvectors, and we describe the spherical harmonic tensor basis in detail in Supplement 6.
@@
Finally, we express our data in terms of the Stokes parameters, a set of four real-valued parameters that are physically interpretable as the intensities measured behind various polarizing filters. The Stokes parameters can also be interpreted as the coefficients of the electric field’s second-moment matrix expanded onto the Pauli matrices (Figure 1e, Supplement 7).
@@ 3.5 Linear contrast-separable shift-invariant imaging systems
Under approximations that are applicable to a wide range of microscopes (details in Section 5 and Supplement 8), we can model our imaging system by splitting our material properties, data channels, and forward model into independent groups called contrast modes, e.g. one label-free mode and several fluorescence modes. Each of these contrast modes is approximately linear across object properties, and each property-to-channel mapping is approximately spatially linear and shift invariant. Therefore, we can express our imaging model as
@@ 3.6 Summary of transfer functions
Figure 2: Key components of the forward model. (a) A dipole emitter (blue arrow) emits polarized fields (red arrows) in an anisotropic pattern (black line, radius proportional to power). (b) Arbitrary linear, circular, and elliptical dipole emissions can be modeled with the Green’s tensor spectrum $G_{ij}(\boldsymbol{\nu})$. Each surface shows the quarter-maximum intensity of on-shell spectral field components (rows) $e_{x}$, $e_{y}$, and $e_{z}$ emitted by dipoles oriented along (columns) $\hat{\textbf{x}}$, $\hat{\textbf{y}}$, and $\hat{\textbf{z}}$, with relative phase shown in color (see color rose). Orthogonal slices through the Green’s tensor spectrum and the Green’s tensor are shown in Figure S3. (c–f) 3D support of various transfer functions with phase encoded in color. (c-d) Real label-free vector transfer functions with $\text{NA}_{\text{det}}=0.75$ and circularly polarized illumination that expresses polarization-resolved data (rows) as outputs of filters that modulate material properties (columns) with (c) low-NA illumination $\text{NA}_{\text{ill}}=0.1$ and (d) high-NA illumination $\text{NA}_{\text{ill}}=0.5$. (e) Scalar label-free transfer function with $\text{NA}_{\text{ill}}=0.5$ and $\text{NA}_{\text{det}}=0.75$. (f) Scalar fluorescence transfer function with $\text{NA}_{\text{det}}=0.75$. See also Video 2.
All transfer functions can be expressed as products and autocorrelations of the illumination, scattering, and detection models, see Table 1 and Supplement 8. We refer to the complete transfer functions as vector models because they account for the complete vectorial nature of light and dipole scattering. We also include scalar models that ignore vector effects, which are reasonable approximations when unpolarized illumination and unpolarized detection are used on isotropic samples.
@@ 4.1 Label-free simulated and real test specimen
Figure 3: Demonstration of vectorial multi-channel reconstruction with experiment and simulation. We imaged (a) (i) a laser-etched spoke pattern of transverse radially anisotropic bubbles. (ii) We made volumetric measurements of the Stokes parameters (columns), and applied a multi-channel label-free reconstruction to (iii) recover material properties. (b) We (i) simulated the phantom’s material properties, (ii) simulated the imaging process, then (iii) simulated the reconstruction. (c) We encoded transverse birefringence properties into color, where brightness indicates the strength of the anisotropy and hue indicate the slow-axis orientation. We compared the orientation and phase reconstructions (rows) in ray- and wave-optics reconstructions (columns). (d) For each data and material property, we measured a series of azimuthal profiles at different radii on the spoke pattern and used the 10th–90th percentile modulation as an empirical estimate of the transverse modulation transfer function (MTF). (e) Similarly, we measured axial profiles through each data and material property. See also Video 3.
@@ 4.2 Cells and tissues
Figure 4: Reconstructions improve visualization and interpretation across length scales in (a) A549 cells, (b) zebrafish, and (c–e) zebrafish regions of interest. (i) Orientation data (left) and reconstructions (right). (ii) Brightfield data (left) and phase reconstructions (right). (iii) Fluorescence data (left) and reconstructions (right). See also Videos 4 and 5.
@@ 4.3 Choice of regularization parameter
waveOrder’s physics-informed reconstructions rely on a list of fixed parameters that specify imaging conditions (e.g., numerical apertures, wavelength, voxel sizes) and a single regularization parameter $\eta$, which penalizes large solutions and suppresses noise. Though [35, 36], in practice, it is common to inspect a regularization sweep and choose an empirical solution (for example, Figure 5 and Figure S1). Regularization parameters that are too small result in noisy reconstructions, while regularization parameters that are too large result in blurry reconstructions. Additionally, we find that the transverse mid-band power, the power in transverse spatial frequencies between one-eighth and one-quarter of the cutoff, is a useful scalar metric that can indicate high-quality reconstructions—see the plateau in Figure 5(d).
@@ 5 Limitations
  • shift invariance, which excludes shift-variant point spread functions [37],
  • weak, single scattering, which excludes thick multiply scattering samples [38], e.g. older zebrafish and thick tissue slices,
@@
  • contrast-separable imaging, which excludes crosstalk between label-free and fluorescence contrast modes [39].
@@ 8 Acknowledgements
We thank Shashvat Mehta for developing the animation shown in Video 1. We thank the Scientific Computing Platform at CZ Biohub for enabling high-performance reconstructions.